determine the integral of function x^2 -x -6/ x +2 dx


1
Expert's answer
2021-08-12T15:29:48-0400

Solve "\int" ( "x^2 - x - \frac{6}{x} +2" ) "dx"


using the properties of integral we can write given integral as,


"\implies" "\int" "x^2" "dx" "-" "\int" "x" "dx" "-" "\int""\frac{6}{x}" "dx" + "\int" 2"dx"


"\implies" "\frac{x^3}{3} - \frac{x^2}{2} - 6ln|x| + 2x + C"


where "C" is an integral constant.